17 December 2009. Last night of SDT LUCIFER observing.

At sunset there is light cirrus. This will affect our flats, but we will take some at J and K anyway. Note, however, that the flats are stable and for whatever programs we may be able to observe, flats from some weeks back could be used.

17:30 - Around this time we obtained sets of K and J flats.

17:40 Wayne reset the encoders after the flat

17:50 Going to correct pointing on ACT0371 (BrG & K in place to avoid saturation). Original CA/IE = -85/-3 New CA/IE = -23/0 We are at elevation 50 deg

18:00 Start Thompson imaging program. Slew to target and execute H script. Lot of coma ... This program makes a large offset to find sky off the galaxy. No guiding or collimation at the sky position. It appears that it thinks it has reached collimation (WFSE ~ 470nm) before making the offset to sky. What we see, however, is an obviously comatic star. When the offset is back to the field, the script does not wait for the collimation. Question of whether collimation was poor because of vignetting or poor collimation. We put guide probe on-axis to collimate - problem was not vignetting. The WFE is increasing, not decreasing. Wayne is going to clear active optics. OK now - we have collimated and seeing is good.

NOTE: for dome seeing, the WFE criterion was increased from 400nm to 600nm. For good nights like tonight (0.5" at start) we question whether we were losing something in image quality. We called John - he did not think this was degrading image quality. The beginning of night issue with strong coma is curious.

18:27 Starting H band imaging script, OSUThompson.scriptH. elev=48deg. Before first image taken after returning from the large offset to sky, the script does not wait for collimation. So this first image may have poorer image quality. Also note that the sky position is not guided.

We can ID the area where the SN should be, however we do not see it in the H band single 33sec sky-subtracted image.

Image FHWM about 0.5"

19:04 Starting K band: OSUThompson.scriptK. Elevation = 41 deg at end.

19:22 Starting Fan's project. Slew to offset/reference star. We identified it. It is at 1062, 1038. we want it at 1022, 1024. Offset -4.75, -1.61. Image it through the slit (104-107). 10-sec through-slit image (108). It is well centered. Then we press return so it is doing a blind offset.

Before starting the observe script, we moved off -20" in Y (relative offset - so down along the slit by 20/0.24" = 83.3 pixels. Spectrum at 0 0 should appear at row 940.67). We did this because we had centered the reference object on row 1024, but realized that the offsets in the observe script would be taking spectra at offset 0 0, so back on row 1024.

Seeing is 0.52" in WFS. First image is 110.

We saw no continuum source at the expected position on the individual sky-subtracted spectra, however we do see something near the top of the slit. This does not appear on the finding charts or an IR (2MASS?) image we pulled up from skycat.

20:53 Start second set at number 0130. Alt = 36.58, T=4.6 and stable, RH=10%. Seeing from WFS is 0.56" and guiding is stable. Note that relative offsets of script end with the spectrum 5" above the origin (0,0) position. So this next set of data will be offset by 5" up.

At 21:55 we are down to 25 deg elevation and seeing ~1" on WFS. The sequence is on 18 of 20 - nearly done.

22:03 Start arc images and FF lamp

22:42 Observing telluric standard. 25 deg elevation. We had to change the star to HIP11766 because the other star was already too low (<10deg) and beyond the collimation limit of the telescope. Taking this standard (nevertheless) in track mode, because pointing was off and the guide star which we had selected was not found. We made a rough pointing correction on this standard, centered it in the slit and observed.

23:00 Starting Woodward's program: We needed to correct pointing. Elev = 49 deg. New CA/IE = -11/-13.

23:30 Starting K band science spectrum. Done - did arc and flat calibrations. We lost our position, so had to re-acquire. In running this acquisition script (which had worked earlier) we had a problem with the Detector focus unit. Walter reinitialized it. We ran the script and again had the error. Walter discussing this with SW group but we are going ahead with the acquisition. We run the science script - but have an error - We commented out the instrument setup and reran. We had a fault initially, and while it was re-initialized, there is a flag which is not being reset and causing errors in the instrument setup.

LST=04:57:51 start K-band spectra, fwhm=0.75", Alt=47.08, humidity=20%, script finished UT=06:30, repeat K-band script a second time at UT=06:37:59/LST=05:02:42

UT = 06:29. Start SpectroPhotometric observation at J.

This is the first SDT SpectroPhot observation with the 3-slit.

Seeing 0.4" at UT=8:10.

LST=06:45 execute K band (image numbers near 216) GD71 spectrophotometric calibration after completing the J-band observations. FWHM=0.34", Alt=69.95. GD71 star white dwarf DA.

UT=08:31, LST=06:56. Take calibrations at this position - for Spectrophot standard.

We are still having this problem with the focus detector unit. Apparently there was an error in the focus initially, but this was fixed. However a flag was not reset - and so all subsequent instrument setups sent from scripts have crashed. Our workaround has been to edited the scripts to comment out the INSTRUMENT SETUP section, do the instrument setup manually and then run this edited script.

08:54 Going to Bechtold program. We tried to manually configure the instrument, but had an error with the camera, filter wheels, grating and flexure - red backgrounds. Nancy clicked initialize - all recovered but Grating Unit. Volker did something (restarted the Instrument Manager) and we are able to run the Bechtold J imaging script and setup the instrument.

09:08 Starting J band imaging (Bechtold). Seeing FWHM 0.37" on WFS. On the 8th of 10 jitters, GCS froze. It had to be killed and restarted and the script killed. Wayne restarted GCS, and we restarted the script - to do 4 more jitters. The collimation for some reason went way off. GCS comes up with first WFE = 0.0. This should not be.

09:56 K band imaging. Counts OK, but probe is in field and vignetting part of field. At K the probe is quite bright and there appears to be some reflection from it - near the probe itself and diagonally opposite it.

Elev = 73 deg. Object is rising.

10:03 We are going to do a phot standard, FS125, for the 2 imaging programs done - but caution - the night is not photometric - there was cirrus early on and may still be some. Standard is saturated in even 2 sec; it has K = 10.355, poor choice--- we need to find one of the fainter standards with K~14. Abandoning this and going to a new program while we look for a new standard.

10:43 On McLinden's program - during acquisition, when trying to move mask to turnout position, we got the MOS problem - and the MOS state went to 'unknown'. Walter, Nancy and Peter looking into it. At 11:40 back on to observing.

11:40 McLinden acquisition of target 1. We found continuum source and centered it on slit. We moved down along slit by 20" so as to not take any spectrum on row 1024. This final acquisition image is #287. We started the science observation script, but then got a MOS error at 12:07. Although the slit was in position, for some unknown reason upon loading the science script, the mask was commanded to move.

12:11 Starting McLinden's target1. #289 is first image. Elevation = 57.6 deg. to 55 deg. Arc done immediately afterwards (we commented out INSTRUMENT SETUP to avoid possibly moving the mask). Calibration flats with HALO1 have far too many counts - not useful. HALO2 flats have counts in the correct range, but the first set (...316...) will have latency from the HALO1. We are repeating only the HALO1 flats.

13:00 Acquiring & observing standard #2 for McLinden's program. Elevation is 57 deg.

13:10 Doing a photometric standard, FS19. Elevation = 42 deg #328...

LST=12:03 had to stop the H-band observations sky background to bright, so move to K-band and continue standard observations (sky 20,000 cts)

LST = 12:15 start Ks sky flats at dawn finish at LST 12:29, blank off Lucifer and head to sleep after starting darks. Nancy monitored the dark script and said that they completed OK (a set for each DIT,NDIT,READMODE and SAVEMODE) however the 180s x 1 darks were contaminated by latency from the overexposed HALO1 lamp flats. She said she'd repeat it later today.

Little cirrus evident in the sky at dawn.

Good night in general.

NOTES:

Observers must provide image scale (e.g., arcsec scale bar or indication) on all finding charts and proper compass rose.

-- OlgaKuhn - 17 Dec 2009
Topic revision: r7 - 18 Dec 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback